OpenStudy (anonymous):

what is the DOMAIN of : f(x) = (-cosx) / 2(rtsqr(1-sinx))

OpenStudy (anonymous):

\[f(x) = \frac{-cosx}{2 \sqrt{1 - sinx}}\] Here the terms inside the square toot must not be negative because square root of negative is not defined so: \[\large 1- sinx \ge 0\] Or: \[sinx \le 1\] Here x can take values for that sinx will not less than or equal to one this is the only condition here..

OpenStudy (anonymous):

just for completing waterineyes statement : \(\sin x≤1\) is true always...but ..we should consider the division by zero...so we must have \(\sin x≠1\) so the domain is \[R-\left\{ 2n\pi+\frac{\pi}{2} \right\} \ \ n=0,1,2,3,...\]

OpenStudy (anonymous):

wait what happens to the cos and 2?